# Charming Darwin: the evolution of QCD parameters across different species

### v1

@@ Abstract
We explore the prospects of measurements of spectral moments of inclusive charm decays with BESIII. The rich and uniquely clean data set of charm mesons and baryons at BESIII offers a unique laboratory to study the evolution of Heavy Quark Expansion (HQE) parameters across different charm hadron species and to shed light on the interplay between heavy quark dynamics and light quark effects. The HQE in terms of inverse powers of the heavy meson mass is well-established in beauty decays, however due to the lighter charm mass its applicability to charm remains an open question. To date no determination of the HQE parameters, the kinetic energy, chromomagnetic moment and Darwin terms, has been attempted. A particular important role here is given to the Darwin and weak annihilation operators, whose values are important to predict lifetimes of heavy hadrons. Using a fast simulation for the BESIII detector response and predictions for spectral moments, we investigate the sensitivity to HQE parameters with today’s and possible future data sets. In addition, we discuss the theory challenges for the HQE in charm and the experimental limitations. We further investigate the sensitivity of determining the CKM matrix element $|V_{cs}|$ with inclusive semileptonic charm decays.
@@ 1 Introduction
The Heavy Quark Expansion (HQE) is a powerful technique to study the dynamics of inclusive heavy hadron decays and predict their lifetimes. In this formalism, the decay rates of heavy quarks can be related to the forward scattering amplitude and the time ordered products can be expanded using the Operator Product Expansion (OPE). The first non-trivial contribution to the decay rate arises from dimension-5 operators (at order $1/m_{Q}^{2}$), where $m_{Q}$ is the mass of the heavy quark and the non-perturbative matrix elements $\{\mu_{\pi}^{2},\mu_{G}^{2}\}$ correspond the kinetic energy and the chromomagnetic moment of the heavy quark inside the heavy meson. Dimension-6 operators (entering at order $1/m_{Q}^{3}$) result in two additional contributions $\{\rho_{D}^{3},\rho_{\mathrm{LS}}^{3}\}$ corresponding to the Darwin and spin-orbit terms. The Darwin operator provides one of the key corrections to the free heavy quark decay lifetime Lenz:2020oce and the HQE is constructed in such a way that it depends on the flavour of the initial heavy hadron.
@@
For inclusive semileptonic beauty meson decays, the HQE has achieved remarkable precision, with hadronic matrix elements identified up to dimension 8 Mannel:2023yqf and $\alpha_{s}^{3}$ corrections to the total decay rate calculated Fael:2020tow . This progress, combined with experimental precision, enables the extraction of inclusive $|V_{cb}|$ at the percent level (see, e.g., Bernlochner:2022ucr ; Finauri:2023kte and references therein). Furthermore, the possibility of measuring inclusive $B_{s}$ decays using a sum-over-exclusive technique has been studied at the LHC DeCian:2023ezb .
@@
For charm decays, the HQE is expected to converge more slowly due to the much smaller expansion scale $\Lambda_{\rm QCD}/m_{c}$. Its validity has been debated over the years Voloshin:2002je ; Falk:1995kn ; Lenz:2013aua ; Lenz:2014jha , with a recent reevaluation in Fael:2019umf . Precisely due to the slower convergence, inclusive charm decays are significantly more sensitive to HQE hadronic matrix elements, such as $\rho_{D}^{3}$, as well as to weak annihilation (WA) operators. These WA operators are crucial input parameters for Standard Model predictions of $B\to X_{u}\ell\bar{\nu}_{\ell}$  Ligeti:2010vd ; Becirevic:2008us ; Gambino:2010jz ; Bigi:2009ym and $B\to X_{s,d}\ell\ell$ Huber:2015sra ; Huber:2019iqf ; Huber:2020vup decays, as well as for lifetime predictions Albrecht:2024oyn . Charm meson and baryon lifetimes have recently been revisited and measured with thus far unprecedented precision by Belle II Belle-II:2021cxx ; Belle-II:2022ggx ; Belle-II:2023eii ; Belle-II:2022plj .
@@
In this paper, we explore the opportunities that the first set of measurements of inclusive semileptonic charm decays could provide. We assess the potential of the large data set recorded by the Beijing Electron Spectrometer Mk. III (BESIII), which is particularly intriguing due to its clean samples of $D^{0}$, $D^{+}$, $D_{s}^{+}$, and $\Lambda_{c}^{+}$ decays. Such an experimental program, combined with corresponding theoretical progress, would a permit a detailed study of the evolution of non-perturbative QCD parameters across various species of heavy hadrons, including the first determination of these parameters in heavy baryon decays. In addition, this program would, for the first time, link semileptonic charm hadron measurements with precision tests of the Standard Model through charm hadron lifetimes.
@@ 2 Prospects for inclusive semileptonic measurements at BESIII
BESIII records electron-positron collisions provided by the Beijing Electron-Positron Collider Mk. II (BEPCII). BEPCII can deliver center-of-mass energies between 1.85 and 4.95 GeV and the BESIII experiment’s large data sets at open-charm-hadron pair-production thresholds provide ideal data samples to perform inclusive semileptonic measurements in the charm system. From the interior to the exterior, the BESIII detector BESIIIDetector is comprised of a drift chamber tracking system, a plastic scintillator time-of-flight system, a crystal calorimeter, a 1T super-conductor solenoid, and a resistive-plate-chamber muon system. The drift chamber provides sub-percent level momentum resolution of charged particles and, combined with measurements from the time-of-flight system, is able to distinguish well between protons, charged pions, and charged kaons. The BESIII calorimeter allows for excellent identification of electrons and provides percent-level resolution on energy depositions of charged and neutral particles.
@@
The sensitivities to HQE parameters from measurements of $D^{+}\to Xe^{+}\nu_{e}$, $D^{0}\to Xe^{+}\nu_{e}$, $D_{s}^{+}\to Xe^{+}\nu_{e}$, and $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ at BESIII are estimated by analyzing Monte Carlo simulation produced with the EvtGen package EvtGen . A measurement strategy similar to that employed in inclusive semileptonic beauty decays at Belle and Belle II BelleQ2 ; BelleIIQ2 is assumed, in which the lepton and all hadrons accompanying the lepton in the final state of the semileptonic decay are reconstructed. This differs from techniques applied in previous measurements of inclusive semileptonic charm decays CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , in which the hadronic system $X$ was not reconstructed and only the charged lepton was identified. These measurements produced precise determinations of the inclusive branching fractions and the electron momentum spectra in the laboratory frame. With additional assumptions on the production mechanism of charmed-hadron pairs, analysis of the measured laboratory-frame distributions allowed for HQE analysis of the $D^{0}$, $D^{+}$, and $D_{s}^{+}$ decays to constrain the contribution of weak-annihilation operators Gambino:2010jz ; Ligeti:2010vd . In this study, we examine how precisely the explicit reconstruction of the hadronic $X$ system can be performed to construct Lorentz-invariant observables. Such quantities are crucial for mitigating the proliferation of HQE parameters at higher orders Fael:2018vsp ; Mannel:2018mqv . This reconstruction enables the reconstruction of the Lorentz-invariant lepton-neutrino mass squared, $q^{2}$, and the Lorentz-invariant mass of the hadronic system, $M_{X}$. Additionally, properties of the $X$ system allow for the separation of $c\to s$ and $c\to d$ transitions, facilitating the first extraction of the CKM matrix elements $|V_{cs}|$ and $|V_{cd}|$ from inclusive semileptonic charm decays.
@@
In the explored analysis procedure, double-tag techniques similar to those developed by the MARKIII collaboration MARKIII_one ; MARKIII_Two are employed with the tag decay modes [^1] Charge conjugation is implied here and throughout the paper. $\overline{D}{}^{0}\to K^{-}\pi^{+}$, $D^{-}\to K^{-}\pi^{+}\pi^{+}$, from Ref. BESIII:2023cym for $D_{s}^{-}$, and from Ref. BESIII:2022cmg for $\Lambda_{c}^{-}$. The dataset listed for BESIII analysis of $D_{s}^{+}$ mesons primarily contains $D_{s}^{+}$ mesons produced through $e^{+}e^{-}\to D_{s}^{*+}D_{s}^{-}$, and so the reconstruction of a $D_{s}^{*}$ meson through a $D_{s}^{*}\to\gamma D_{s}$ is also assumed. The assumed size of BESIII data samples are listed Table 1. The listed data sets correspond to samples which BESIII has analyzed for publication, with the exception of the samples for $D^{0}$ and $D^{+}$ analysis, where a larger dataset has recently been collected.
@@ 2.1 Estimated distributions from fast simulation
A fast-simulation software based on EvtGen EvtGen is developed to estimate the effects of BESIII detector response, momentum resolution, and geometric and kinematic acceptance on inclusive semileptonic measurements. The effects of final-state radiation are neglected in the simulation. The decays of charmed hadrons are generated by EvtGen for the tag-side meson and all observed semileptonic decays PDG2022 , and those predicted by isotopic symmetry based on observed decays. The ISGW2 ISGW2 model is assumed for all $\{D^{+},D^{0},D_{s}^{+},\Lambda_{c}^{+}\}\to Me^{+}\nu$ decays, where $M$ is any hadron, and a uniform phase-space model is assumed when more than one hadron is produced in the semileptonic decay. Detector effects are incorporated probabilistically based on publicly reported reconstruction efficiencies and resolutions. The efficiencies of tracking charged particles and particle identification are estimated based on the published measurement of the branching fraction of $D_{s}^{+}\to Xe^{+}\nu_{e}$ decays BESIIIDsIncSL . Photon reconstruction efficiencies and resolution are estimated based on Ref. BESIIINeutrals . The energy deposition of $K^{0}_{L}$ mesons and neutrons in the BESIII calorimeter is estimated based on the analysis of $K^{0}_{L}$ backgrounds in the search for $D^{0}\to\pi^{0}\nu\overline{\nu}$ at BESIII BESIIIPiNuNu . The reconstruction efficiency and momentum resolution of $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays is estimated based on Ref. BESIIIIncLKS . The efficiency and momentum resolution of reconstructing $\Lambda\to p\pi^{-}$ decays is assumed to be identical to $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays due to the similar vertex reconstruction.
@@
Simulated measured distributions of the kinematic distributions of interest in the HQE (the squared four-momentum of the $e^{+}\nu$ system $q_{\text{reco}}^{2}$, the $e^{+}$ energy in the parent rest frame $E_{\ell}$, and the mass of the hadronic system in the final state $M_{X}$) are produced by the fast simulation based on the estimated double-tag sample sizes listed in Table 1. The BESIII detector is unable to consistently distinguish between electrons and other charged particles for candidate tracks with momenta less than 200 MeV. As such, it is required that the momentum of the electron candidate be above this threshold. The resulting $E_{\text{miss}}-p_{\text{miss}}$ distribution produced by the fast simulation of for each charmed hadron is shown in Fig. 1, with the contribution of decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays in the final state highlighted. As the figure shows, decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ are often poorly reconstructed, due to the poor resolution of $K^{0}_{L}$ energy depositions in the calorimeter and the inability of the BESIII detector to distinguish between energy depositions due to photons and those due to $K^{0}_{L}$ mesons. As such, a requirement of $\left|E_{\text{miss}}-p_{\text{miss}}\right|<500\text{ MeV}$ is placed on simulated events to estimate the sensitivity of measurements at BESIII, and a correction is later applied to account for this requirement. It should be noted that a measurement at BESIII could correct for decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays in the final state that are removed with this selection based on measurements of decays with $K^{0}_{S}\to\pi^{+}\pi^{-}$ decays in the final state, which BESIII reconstructs with high efficiency and good momentum resolution. Contributions from neutrons, primarily in $\Lambda_{c}^{+}$ decays, can similarly be corrected for assuming isotopic symmetry [^2] The decay $\Lambda_{c}^{+}\to ne^{+}\nu$ must be handled separately. based on observed decays with protons in the final state.
@@
The kinematic and geometric acceptance of BESIII associated with reconstructing and identifying the positron and the $E_{\text{miss}}-p_{\text{miss}}$ requirement reduce the sample size for analysis, and sculpt the kinematic distributions of interest. The efficiency of by the acceptance and selections is approximately 68%, 53%, 44%, and 49% for the $D^{0}\to Xe^{+}\nu_{e}$, $D^{+}\to Xe^{+}\nu_{e}$, $D_{s}^{+}\to Xe^{+}\nu_{e}$, and $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ samples, respectively. While the resolution is predicted to be considerably worse for the $D_{s}^{+}$ sample, a measurement at BESIII could improve this with a kinematic constraint on the $D_{s}^{*}$ candidate. Comparisons between the generated kinematic distributions of interest of $D^{+}\to Xe^{+}\nu_{e}$ decays and the corresponding reconstructed distributions produced by the fast simulation are shown in Fig. 2.
@@
Figure 4: Estimated $M_{X}$ distribution of (left) $c\to s$-tagged $D^{+}$ semileptonic decays and (right) $c\to d$-tagged $D^{+}$ semileptonic decays. In each case, the expected observed distribution is shown in light blue, the contribution of events with $K^{0}_{L}$ mesons and $K_{S}^{0}\to\pi^{0}\pi^{0}$ decays in the final state is shown in dark blue, and the true distribution of events passing selections shown as the dashed orange histogram.

### v3

@@ Abstract
We explore the prospects of measurements of spectral moments of inclusive charm decays with BESIII. The rich and uniquely clean data set of charm mesons and baryons at BESIII offers a unique laboratory to study the evolution of Heavy Quark Expansion (HQE) parameters across different charm hadron species and to shed light on the interplay between heavy quark dynamics and light quark effects. The HQE in terms of inverse powers of the heavy meson mass is well-established in beauty decays, however due to the lighter charm mass its applicability to charm remains an open question. To date no determination of the HQE parameters, the kinetic energy, chromomagnetic moment and Darwin terms, has been attempted. A particularly important role here is given to the Darwin and weak annihilation operators, whose values are important to predict lifetimes of heavy hadrons. Using a fast simulation for the BESIII detector response and predictions for spectral moments, we investigate the sensitivity to HQE parameters with today’s and possible future data sets. In addition, we discuss the theory challenges for the HQE in charm and the experimental limitations. We further investigate the sensitivity of determining the CKM matrix element $|V_{cs}|$ with inclusive semileptonic charm decays.
@@ 1 Introduction
The Heavy Quark Expansion (HQE) is a powerful technique to study the dynamics of inclusive heavy hadron decays and predict their lifetimes. In this formalism, the decay rates of heavy quarks can be related to the forward scattering amplitude and the time ordered products can be expanded using the Operator Product Expansion (OPE). The first non-trivial contribution to the decay rate arises from dimension-5 operators (at order $1/m_{Q}^{2}$), where $m_{Q}$ is the mass of the heavy quark and the non-perturbative matrix elements $\{\mu_{\pi}^{2},\mu_{G}^{2}\}$ correspond to the kinetic energy and the chromomagnetic moment of the heavy quark inside the heavy meson. Dimension-6 operators (entering at order $1/m_{Q}^{3}$) result in two additional contributions $\{\rho_{D}^{3},\rho_{\mathrm{LS}}^{3}\}$ corresponding to the Darwin and spin-orbit terms. The Darwin operator provides one of the key corrections to the free heavy quark decay lifetime Lenz:2020oce and the HQE is constructed in such a way that it depends on the flavour of the initial heavy hadron.
@@
For inclusive semileptonic beauty meson decays, the HQE has achieved remarkable precision, with hadronic matrix elements identified up to dimension 8 Mannel:2023yqf and $\alpha_{s}^{3}$ corrections to the total decay rate calculated Fael:2020tow , collected in the open-source package Kolya Fael:2024fkt . This progress, combined with experimental precision, enables the extraction of inclusive $|V_{cb}|$ at the percent level (see, e.g., Bernlochner:2022ucr ; Finauri:2023kte and references therein). Furthermore, the possibility of measuring inclusive $B_{s}$ decays using a sum-over-exclusive technique has been studied at the LHC DeCian:2023ezb .
For charm decays, the challenge is the low mass of the charm which lies close to the non-perturbative QCD regime, which may lead to a break down of the OPE. For example, four-quark operators, weak annihilation, and Pauli interference terms enter in the HQE for charm meson lifetimes. These terms are numerically enhanced by $16\pi^{2}$ and play an important role in explaining the large lifetime differences in charm decays Lenz:2013aua ; Lenz:2014jha ; Belle-II:2021cxx .
@@
On the other hand, the semileptonic widths of the different charm decays are more similar, which may indicate that weak annihilation (WA) operators are of less importance. This gives confidence in the validity of the OPE, which was discussed in  Voloshin:2002je ; Falk:1995kn , with a recent reevaluation in Fael:2019umf . Precisely due to the slower convergence, inclusive charm decays are significantly more sensitive to HQE hadronic matrix elements, such as $\rho_{D}^{3}$, as well as to WA operators. These WA operators are crucial input parameters for Standard Model predictions of $B\to X_{u}\ell\bar{\nu}_{\ell}$  Ligeti:2010vd ; Becirevic:2008us ; Gambino:2010jz ; Bigi:2009ym and $B\to X_{s,d}\ell\ell$ Huber:2015sra ; Huber:2019iqf ; Huber:2020vup decays. As discussed above they also form crucial inputs in lifetime predictions Albrecht:2024oyn . Charm meson and baryon lifetimes have recently been revisited and measured with thus far unprecedented precision by Belle II Belle-II:2021cxx ; Belle-II:2022ggx ; Belle-II:2023eii ; Belle-II:2022plj .
@@
In this paper, we explore the opportunities that measurements of inclusive semileptonic charm decays could provide. We assess the potential of the large data set recorded by the Beijing Electron Spectrometer Mk. III (BESIII), which is particularly intriguing due to its clean samples of $D^{0}$, $D^{+}$, $D_{s}^{+}$, and $\Lambda_{c}^{+}$ decays. Such an experimental program, combined with corresponding theoretical progress, would permit a detailed study of the evolution of non-perturbative QCD parameters across various species of heavy hadrons, including the first determination of these parameters in heavy baryon decays. In addition, this program would, for the first time, link semileptonic charm hadron measurements with precision tests of the Standard Model through charm hadron lifetimes.
@@ 2 Prospects for inclusive semileptonic measurements at BESIII
BESIII records electron-positron collisions provided by the Beijing Electron-Positron Collider Mk. II (BEPCII) BEPCII . BEPCII can deliver center-of-mass energies between 1.85 and 4.95 GeV and the BESIII experiment’s large data sets at open-charm-hadron pair-production thresholds provide ideal data samples to perform inclusive semileptonic measurements in the charm system. From the interior to the exterior, the BESIII detector BESIIIDetector is comprised of a drift chamber tracking system, a plastic scintillator time-of-flight system, a crystal calorimeter, a 1T super-conductor solenoid, and a resistive-plate-chamber muon system. The drift chamber provides sub-percent level momentum resolution of charged particles and, combined with measurements from the time-of-flight system, is able to distinguish well between protons, charged pions, and charged kaons. The BESIII calorimeter allows for excellent identification of electrons and provides percent-level resolution on energy depositions of charged and neutral particles.
@@
The sensitivities to HQE parameters from measurements of $D^{+}\to Xe^{+}\nu_{e}$, $D^{0}\to Xe^{+}\nu_{e}$, $D_{s}^{+}\to Xe^{+}\nu_{e}$, and $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ at BESIII are estimated by analyzing Monte Carlo simulation produced with the EvtGen package EvtGen . A measurement strategy similar to that employed in inclusive semileptonic beauty decays at Belle and Belle II BelleQ2 ; BelleIIQ2 is assumed, in which the lepton and all hadrons accompanying the lepton in the final state of the semileptonic decay are reconstructed. This differs from techniques applied in previous measurements of inclusive semileptonic charm decays CLEO:2009uah ; BESIIIDsIncSL ; BESIII:2022cmg , in which the hadronic system $X$ was not reconstructed and only the charged lepton was identified. These measurements produced precise determinations of the inclusive branching fractions and the electron momentum spectra in the laboratory frame. With additional assumptions on the production mechanism of charm-hadron pairs, the measured laboratory-frame distributions allowed for HQE analysis of the $D^{0}$, $D^{+}$, and $D_{s}^{+}$ decays to constrain the contribution of weak-annihilation operators Gambino:2010jz ; Ligeti:2010vd . In this study, we examine how precisely the explicit reconstruction of the hadronic $X$ system can be performed to construct Lorentz-invariant observables. Such quantities are crucial for mitigating the proliferation of HQE parameters at higher orders Fael:2018vsp ; Mannel:2018mqv . This reconstruction enables the reconstruction of the Lorentz-invariant lepton-neutrino mass squared, $q^{2}$, and the Lorentz-invariant mass of the hadronic system, $M_{X}$. Additionally, properties of the $X$ system allow for the separation of $c\to s$ and $c\to d$ transitions, facilitating the first extraction of the CKM matrix elements $|V_{cs}|$ and $|V_{cd}|$ from inclusive semileptonic charm decays.
@@
In the explored analysis procedure, double-tag techniques similar to those developed by the MARKIII collaboration MARKIII_one ; MARKIII_Two are employed with the tag decay modes [^1] Charge conjugation is implied here and throughout the paper. $\overline{D}{}^{0}\to K^{-}\pi^{+}$, $D^{-}\to K^{-}\pi^{+}\pi^{+}$, from Ref. BESIII:2023cym for $D_{s}^{-}$, and from Ref. BESIII:2022cmg for $\Lambda_{c}^{-}$. The dataset listed for BESIII analysis of $D_{s}^{+}$ mesons primarily contains $D_{s}^{+}$ mesons produced through $e^{+}e^{-}\to D_{s}^{*\pm}D_{s}^{\mp}$, and so the identification of a $D_{s}^{*}$ meson through a $D_{s}^{*}\to D_{s}\gamma$ is also assumed, where the $D_{s}$ meson produced through the $D_{s}^{*}$ can decay to either a semileptonic decay or to a hadronic tag final-state. The assumed size of BESIII data samples are listed Table 1. The listed data sets correspond to samples which BESIII has analyzed for publication, with the exception of the samples for $D^{0}$ and $D^{+}$ analysis, where a larger dataset has recently been collected.
@@ 2.1 Estimated distributions from fast simulation
A fast-simulation software based on EvtGen EvtGen is developed to estimate the effects of BESIII detector response, momentum resolution, and geometric and kinematic acceptance on inclusive semileptonic measurements. The effects of final-state radiation are neglected in the simulation. The decays of charm hadrons are generated by EvtGen for the tag-side meson and all observed semileptonic decays PDG2022 , and those predicted by isotopic symmetry based on observed decays. The ISGW2 ISGW2 model is assumed for all $\{D^{+},D^{0},D_{s}^{+},\Lambda_{c}^{+}\}\to Me^{+}\nu$ decays, where $M$ is any hadron, and a uniform phase-space model is assumed when more than one hadron is produced in the semileptonic decay. Detector effects are incorporated probabilistically based on publicly reported reconstruction efficiencies and resolutions. The efficiencies of tracking charged particles and particle identification are estimated based on the published measurement of the branching fraction of $D_{s}^{+}\to Xe^{+}\nu_{e}$ decays BESIIIDsIncSL . Photon reconstruction efficiencies and resolution are estimated based on Ref. BESIIINeutrals . The energy deposition of $K^{0}_{L}$ mesons and neutrons in the BESIII calorimeter is estimated based on the analysis of $K^{0}_{L}$ backgrounds in the search for $D^{0}\to\pi^{0}\nu\overline{\nu}$ at BESIII BESIIIPiNuNu . The reconstruction efficiency and momentum resolution of $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays is estimated based on Ref. BESIIIIncLKS . The efficiency and momentum resolution of reconstructing $\Lambda\to p\pi^{-}$ decays is assumed to be identical to $K_{S}^{0}\to\pi^{+}\pi^{-}$ decays due to the similar vertex reconstruction.
@@
Simulated measured distributions of the kinematic distributions of interest in the HQE (the squared four-momentum of the $e^{+}\nu$ system $q_{\text{reco}}^{2}$, the $e^{+}$ energy in the parent rest frame $E_{\ell}$, and the mass of the hadronic system in the final state $M_{X}$) are produced by the fast simulation based on the estimated double-tag sample sizes listed in Table 1. The BESIII detector is unable to consistently distinguish between electrons and other charged particles for candidate tracks with momenta less than 200 MeV. As such, it is required that the momentum of the electron candidate be above this threshold. The resulting $E_{\text{miss}}-p_{\text{miss}}$ distribution produced by the fast simulation for each charm hadron is shown in Fig. 1, with the contribution of decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays in the final state highlighted. As the figure shows, decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays are often poorly reconstructed, due to the poor resolution of $K^{0}_{L}$ energy depositions in the calorimeter and the inability of the BESIII detector to distinguish between energy depositions due to photons and those due to $K^{0}_{L}$ mesons. As such, a requirement of $\left|E_{\text{miss}}-p_{\text{miss}}\right|<500\text{ MeV}$ is placed on simulated events to estimate the sensitivity of measurements at BESIII, and a correction is later applied to account for this requirement. It should be noted that a measurement at BESIII could correct for decays with $K^{0}_{L}$ mesons or $K^{0}_{S}\to\pi^{0}\pi^{0}$ decays in the final state that are removed with this selection based on measurements of decays with $K^{0}_{S}\to\pi^{+}\pi^{-}$ decays in the final state, which BESIII reconstructs with high efficiency and good momentum resolution. Contributions from neutrons, primarily in $\Lambda_{c}^{+}$ decays, can similarly be corrected for assuming isotopic symmetry [^2] The decay $\Lambda_{c}^{+}\to ne^{+}\nu$ must be handled separately. based on observed decays with protons in the final state.
The decay $D_{s}^{+}\to\tau^{+}\nu_{\tau},\tau^{+}\to e^{+}\nu_{e}\overline{\nu}_{\tau}$ is expected to pass the selection requirements at a similar rate to signal processes, as it did in the previous BESIII analysis of $D_{s}^{+}\to Xe^{+}\nu$ BESIIIDsIncSL . We suggest that it can be handled in a similar fashion in our procedure as the previous analysis, subtracting on its contribution based off of the measured branching fraction, as the modeling of this decay carries negligible systematic uncertainty.
@@
The kinematic and geometric acceptance of BESIII associated with reconstructing and identifying the positron and the $E_{\text{miss}}-p_{\text{miss}}$ requirement reduce the sample size for analysis, and sculpt the kinematic distributions of interest. The efficiency of the acceptance and selections is approximately 68%, 53%, 44%, and 49% for the $D^{0}\to Xe^{+}\nu_{e}$, $D^{+}\to Xe^{+}\nu_{e}$, $D_{s}^{+}\to Xe^{+}\nu_{e}$, and $\Lambda_{c}^{+}\to Xe^{+}\nu_{e}$ samples, respectively. While the resolution is predicted to be considerably worse for the $D_{s}^{+}$ sample, a measurement at BESIII could improve this with a kinematic constraint on the $D_{s}^{*}$ candidate. Comparisons between the generated kinematic distributions of interest of $D^{+}\to Xe^{+}\nu_{e}$ decays and the corresponding reconstructed distributions produced by the fast simulation are shown in Fig. 2.
@@
Figure 4: Estimated $M_{X}$ distribution of (left) $c\to s$-tagged $D^{+}$ semileptonic decays and (right) $c\to d$-tagged $D^{+}$ semileptonic decays. In each case, the expected observed distribution is shown in light blue, the contribution of events with $K^{0}_{L}$ mesons and $K_{S}^{0}\to\pi^{0}\pi^{0}$ decays in the final state is shown in dark blue, and the true distribution of events passing selections shown as the dashed green histogram.